OpenStudy (anonymous):

4/x + 1/x-2 =1 Please give me an answer! I need a hero here!! I'm running out of time.

OpenStudy (anonymous):

He's gotta be strong And he's gotta be fast And he's gotta be fresh from the fight!

OpenStudy (anonymous):

\[\frac{4}{x} + \frac{1}{x-2} = 1?\]

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

\[\frac{4x -8}{x(x-2)} + \frac{x}{x(x-2)}\]

OpenStudy (anonymous):

\[\frac{5x-8}{x(x-2)}\]

OpenStudy (anonymous):

4/x + 1/(x-2) = 1 4(x-2)/(x-2)x + x/x(x-2) = 1 [(4x-8)+x]/(x^2 - 2x) = 1 (4x-8)+x = (x^2 - 2x) 5x-8 = x^2 - 2x x^2 - 7x + 8 = 0 (x-8)(x+1)=0 x=8, x=-1

OpenStudy (anonymous):

\[x=\frac{7\pm\sqrt{17}}{2}\]

OpenStudy (anonymous):

*hug* thanks so so so much, wait a sec, if you dont mind. i have more

OpenStudy (anonymous):

go ahead and post

OpenStudy (anonymous):

ok just a sec :)

OpenStudy (anonymous):

\[2/x^2-3x+2=3x/x-1-x/x-2\]

OpenStudy (anonymous):

\[\frac{2}{x^2-3x+2}=\frac{3x}{x-1}-\frac{x}{x-2}\] like that?

OpenStudy (anonymous):

if so try \[x=\frac{5\pm\sqrt{41}}{4}\]

OpenStudy (anonymous):

ok be right back... :)

OpenStudy (anonymous):

awesome thanks so much! can you help me with two more? thats all i have left :D

OpenStudy (anonymous):

\[4x=\sqrt{2x+3}\]